### v1

@@ 1 Introduction
Coral reef surveys utilize various techniques and platforms, including satellite and aerial remote sensing, vessel-based sonar and LiDAR, underwater vehicle-based imaging, and manual in-situ underwater surveys (Collin et al. (2018); Character et al. (2021); Price et al. (2019); Rossi et al. (2020a)). Satellite and aerial methods provide rapid data acquisition for large-scale monitoring (Casella et al. (2017)) but lack the accuracy and sufficient resolution to capture the intricate details of coral reef structures. Manual surveys, on the other hand, are labor-intensive and limited by their spatial and temporal scope. While Sonar and LiDAR enhance the acquisition of geometric data from benthic habitats, their limited spatial resolution and high cost make it difficult to capture fine details. The emergence of vision-based underwater observation techniques has allowed for the collection of high-resolution images from close distances at low cost. This advancement facilitates the use of emerging image-based 3D reconstruction methods to produce high-accuracy and high-resolution 3D models of seabed coral reefs in a cost-effective, non-invasive manner (Rossi et al. (2020a); Zhong et al. (2023b); Lange and Perry (2020)). Furthermore, these techniques support the production of realistic and richly textured orthomosaics, digital surface models. Integrating them with other survey methods across various environments can enhance the accuracy and efficiency of 3D coral habitat reconstruction, leading to more comprehensive and detailed assessments of coral reef ecology and environmental changes.
@@
Early studies made preliminary attempts to use traditional photogrammetric methods for reconstructing underwater coral reefs (Andono et al. (2012); Drap et al. (2013); Guo et al. (2016)). With the advancements in photogrammetric computer vision, the emergency of 3D reconstruction techniques such as Structure-from-Motion (SfM) (Schonberger and Frahm (2016)) and Multi-View Stereo (MVS) (Schönberger et al. (2016)) has enabled the automation of the reconstruction process. The typical photogrammetric 3D reconstruction workflow typically comprises two key stages: camera pose estimation and dense surface reconstruction. Camera pose estimation, achieved through techniques like SfM, determines the position and orientation of the cameras in 3D space by performing tasks like feature extraction, feature matching, and bundle adjustment. Dense surface reconstruction, on the other hand, focuses on generating a detailed model of the scene, often trying to estimate the 3D coordinates of each pixel, with MVS being the current predominant approach. Researchers have applied these techniques to 3D reconstruction of underwater coral reefs, achieving impressive results with 3D measurements accurate to the centimeter or even millimeter level (Zhong et al. (2023b); Kalacska et al. (2018); Mohamed et al. (2020)). However, these methods have limitations and may experience degraded performance or even failure under suboptimal conditions. Specifically, issues arise when images exhibit sparse or chaotic textures, high noise levels, occlusions, or insufficient overlap between images. Given the variable imaging conditions underwater and the complex, intricate structures of coral reefs, such as their tentacles, these challenging scenarios frequently arise, posing substantial demands on the accuracy, robustness, and efficiency of underwater 3D reconstruction technologies.
@@
Despite the impressive performance of these advanced methods in testing environments, their applicability in underwater environments, particularly in coral reef scenes, remains to be further validated. Can these methods genuinely outperform traditional approaches when processing coral reef images? Which method represents the optimal choice for more effective coral reef reconstruction? Addressing these questions is essential for the practical applications. Nevertheless, the lack of thorough reviews and evaluations of these advanced solutions has resulted in a significant gap between technical innovations and practical applications. As coral reef monitoring and conservation become more urgent, there is a growing demand for clear answers and guidance in this field. This motivates us to do this study and write this paper, which systematically summarizes current mainstream image-based 3D reconstruction methods and advanced techniques, evaluates them through extensive experiments, and provides recommendations and discussions about existing issues and potential future research directions.
2 The Workflow of 3D Reconstruction
@@ 2.1 Data collection and preparation
In the first stage, the primary objective is to acquire high-resolution underwater images using visual sensors such as Digital Single Lens Reflex (DSLR) Camera and action cameras. The captured images should be in high definition, overlap with each other, and cover the entire survey area. This is essential for applying photogrammetric computer vision techniques for the 3D reconstruction of coral reefs. In practice, data collection is typically performed by divers or unmanned vehicles such as Remotely Operated Vehicles (ROVs) or Autonomous Underwater Vehicles (AUVs).
@@ 2.2 Camera pose estimation
Camera pose estimation focuses on determining the accurate positions and orientations of the cameras. Traditional aerial triangulation relies on initial estimates of camera poses to achieve accurate results, which can be a limitation in underwater environments that often lack reliable navigation data. In contrast, SfM automatically estimates camera poses without prior information, making it a powerful solution. Its speed, low cost, simplicity, and versatility have led to widespread adoption in 3D reconstruction tasks.
As illustrated in Figure 1, SfM is primarily composed of two main components: correspondence search and reconstruction. The first component aims to detect regions of overlap between input images and match corresponding projections of the same 3D points across overlapping areas. This process involves feature extraction and matching, where features are a combination of keypoints, which are distinct points in an image, and their descriptors, which describe the surrounding area of each keypoint. These features are typically extracted using local feature methods, ensuring they are invariant to changes in translation, rotation, scale, and illumination. Subsequently, the matching process seeks to identify corresponding projections of the same points across overlapping images, thereby establishing correspondences between image pairs. Feature matching commonly involves geometrical verification to eliminate potential outliers in the matches (Zhong et al. (2023c)). This verification is typically based on local photometric and geometric constraints. The quality of the correspondence search directly impacts the accuracy and reliability of the reconstruction. Existing solutions may face significant challenges in complex application scenarios, highlighting this as a key research problem.
The second component aims to estimate the image poses and 3D scene points using the correspondences between image pairs, which is achieved through image registration and triangulation in SfM reconstruction. The crucial processing step in this phase is bundle adjustment (Triggs et al. (2000)). Given that the extracted feature locations and correspondences may contain errors, and lens distortions can occur, errors can accumulate rapidly, potentially leading to drift in SfM. Bundle adjustment addresses this by minimizing the reprojection error—the discrepancy between the projected positions of 3D points in images and the actual locations of detected feature points—thereby improving the accuracy and consistency of the reconstruction. The object function of bundle adjustment is presented as follow:
@@
[…]
2.3 Dense surface reconstruction
Although a 3D point cloud can be generated through triangulation during SfM reconstruction, it is often too sparse to represent the detailed geometric information of the observed area. This limitation necessitates the use of dense surface reconstruction techniques to recover the finer details of the scene. Contemporary methods for dense reconstruction employ a range of approaches (discussed in detail in Section 4), but they all share a common principle of generating a dense representation of the 3D scene, such as a dense point cloud or mesh, by integrating information from multiple views. This process depends on the geometric relationships and viewpoint consistency among multi-view images for inference and optimization. Consequently, it necessitates the intrinsic and extrinsic camera parameters, which can be derived from SfM reconstruction. The quality of the dense surface reconstruction directly impacts subsequent applications such as measurement, scene analysis, and digital modeling.
3 Camera pose estimation solutions
This section provides a brief overview of the technologies involved in camera pose estimation. Given that current SfM solutions have well-established processes for pose estimation, triangulation, and bundle adjustment, the primary challenge is correspondence search. This process can be unreliable, especially in scenes with chaotic or sparse textures or under significant variations in viewpoint or lighting conditions, which can hinder SfM reconstruction. Due to the greater impact of the water medium on imaging compared to the air medium, coral reef images captured underwater face challenges such as image degradation, similarity of natural textures, and cluttered background textures, which make correspondence searching in images significantly more challenging. Figure 2 illustrates a flexible framework for incremental SfM. This section will delve into three technologies within this framework: feature extraction, feature matching, and reconstruction.
@@ 3 Camera pose estimation solutions
3.1 Feature extraction
SfM relies on local features to align overlapping images by identifying the same scene points. These features must be invariant to radiometric and geometric changes to ensure accurate correspondence between different images. Traditional hand-crafted feature extraction methods typically follow a two-stage pipeline: keypoint detection followed by descriptor computing. The geometric invariance of these descriptors primarily addresses scale invariance and orientation invariance, which have been extensively studied over the past decades. SIFT (Lowe (1999, 2004)) is widely known for its scale and orientation invariance by estimating keypoint scales using gradient histograms. SURF (Bay et al. (2006)) significantly improves the computational efficiency of SIFT through various optimizations, while ORB features (Rublee et al. (2011)) further simplify and accelerate the process. Alcantarilla et al. introduced KAZE features (Alcantarilla et al. (2012a)), which leverage non-linear scale space through non-linear diffusion filtering to achieve invariance to rotation, scale, and limited affine transformations. They also proposed an enhanced version, AKAZE (Alcantarilla and Solutions (2011)), which improves processing speed. However, these methods are based on heuristics, which can be difficult to generalize across different scenarios, potentially leading to reconstruction failures.
@@ 3.1 Feature extraction
Hand-crafted methods SIFT 1999 Advantages: (1) Rapid computational performance; (2) Simple and comprehensible principles; (3) High accuracy in keypoint localization. Disadvantages: (1) Significant sensitivity to parameter settings; (2) Performance instability under extreme or challenging conditions.
@@
Deep learning-based methods LIFT 2016 Advantages: (1) Enhanced descriptive capability and robustness; (2) Simple parameter adjustment. Disadvantages: (1) Greater complexity of the model architecture; (2) Dependence on the datasets; (3) Require more computational power.
@@
3.2 Feature matching
After feature extraction, a set of keypoints and their corresponding descriptors are obtained. For SfM reconstruction, accurate feature matching is required to establish correspondences between images of the same scene region. The problem is classically solved by matching a keypoint with its most similar counterpart in another image, specifically identifying the nearest neighbor in descriptor space. However, under conditions such as low image overlap, sparse or overly similar textures, and variations in lighting, descriptors may fail to accurately represent features. This can result in many outliers—matches that are incorrect or significantly different from the expected correspondences. The classical ratio-test (Lowe (2004)) s a basic and straightforward feature matching method, often used in conjunction with mutual nearest neighbor checks for rapid correspondence search. While these simple filters are efficient and widely applied, their performance is limited. They often miss many outliers or incorrectly filter out inliers, which can complicate SfM solutions. To achieve more robust and reliable feature matching, researchers have explored techniques such as local spatial consistency checks and global geometric verification. Among these methods, one of the most representative is AdaLAM (Cavalli et al. (2020)), proposed by Cavalli et al. AdaLAM demonstrates significant robustness and is notable for its ease of deployment and high operational efficiency. While AdaLAM integrates several ideas from previous research, it remains based on geometrical assumptions, such as local affine consistency, which may limit its applicability in certain scenarios—especially when dealing with no-planar surfaces. To address this, it employs adaptive strategies to relax these assumptions, enhancing its generalization across different domains. Compared to general methods, AdaLAM not only significantly reduces outliers but also increases the number of inliers.
@@ 3.3 SfM reconstruction
4 Dense surface reconstruction solutions
Dense surface reconstruction generates a detailed, dense sampling of the surface geometry of coral reefs, enabling fine-grained measurement and analysis. However, the intricate and varied morphology of coral reefs, such as their small tentacles, presents challenges including occlusions, shadows and similar structures. Additionally, factors such as lighting, water conditions, and sediment may affect the accuracy of the reconstruction. Moreover, since the structural changes in coral reefs over a year can be merely on the order of centimeters, achieving high precision in reconstruction is essential. Therefore, there is an urgent need for high-quality and efficient dense reconstruction solutions. This section focuses on the current dense surface reconstruction technologies applicable to coral reefs. As illustrated in Figure 3, these methods can be broadly categorized into four categories: traditional multi-view stereo (MVS) methods, deep learning-based MVS methods, methods based on neural radiance fields (NeRFs), and methods based on Gaussian Splatting (GS). Representative approaches within these categories are summarized in Table 2 and will be discussed in detail in the following sections.

### v2

@@ 1 Introduction
Coral reef surveys are conducted using a variety of platforms, including satellites, aerial systems, vessels, underwater vehicles, and manual in-situ methods (Collin et al. (2018); Price et al. (2019); Rossi et al. (2020); Costa et al. (2009)). Manual surveys are labor-intensive and limited by their spatial and temporal scope. Satellite and aerial platforms equipped with cameras or LiDAR systems enable large-scale and frequent monitoring (Casella et al. (2017); Collin et al. (2018)). However, due to their limited spatial resolution and susceptibility to surface water interference, they are generally insufficient for capturing the fine-scale structural complexity of coral reefs. Vessel-based approaches typically employ sonar, which is effective for wide-area mapping but less suited for shallow waters and incapable of capturing spectral or color information (Costa et al. (2009)). Underwater LiDAR, which is capable of close-range, high-precision surveying of coral formations (Vogler et al. (2019)), faces challenges including high costs, operational complexity, and limited capacity for capturing color information. The emergence of vision-based underwater observation techniques has allowed for the collection of high-resolution images from close distances at low cost. This advancement facilitates the use of emerging image-based 3D reconstruction methods to produce high-accuracy and high-resolution 3D models of seabed coral reefs in a cost-effective, non-invasive manner (Rossi et al. (2020); Zhong et al. (2023b); Lange and Perry (2020)). Furthermore, these techniques support the production of realistic and richly textured orthomosaics, digital surface models. Integrating them with other survey methods across various environments can enhance the accuracy and efficiency of 3D coral habitat reconstruction, leading to more comprehensive and detailed assessments of coral reef ecology and environmental changes.
@@
Early studies made preliminary attempts to use traditional photogrammetric methods for reconstructing underwater coral reefs (Andono et al. (2012); Drap et al. (2013); Guo et al. (2016)). Some research further incorporated photogrammetry with an underwater geodetic network to establish a consistent and unambiguous reference frame, thereby enabling precise coral monitoring (Nocerino et al. (2020); Carlot et al. (2020); Neyer et al. (2018); Rossi et al. (2020); Casella et al. (2017)). With the advancements in photogrammetric computer vision, the emergence of image-based 3D reconstruction techniques such as Structure-from-Motion (SfM) (Schonberger and Frahm (2016)) and Multi-View Stereo (MVS) (Schönberger et al. (2016)) has enabled the automation of the reconstruction process. The typical photogrammetric 3D reconstruction workflow comprises two key stages: camera pose estimation and dense surface reconstruction. Camera pose estimation, achieved through techniques like SfM, determines the position and orientation of the cameras in 3D space by performing tasks like feature extraction, feature matching, and bundle adjustment. Dense surface reconstruction, on the other hand, focuses on generating a detailed model of the scene, often trying to estimate the 3D coordinates of each pixel, with MVS being the current predominant approach. Researchers have applied these techniques to 3D reconstruction of underwater coral reefs, achieving impressive results with 3D measurements accurate to the centimeter or even millimeter level (Zhong et al. (2023b); Kalacska et al. (2018); Mohamed et al. (2020)). However, these methods have limitations and may experience degraded performance or even failure under suboptimal conditions. Specifically, issues arise when images exhibit sparse or chaotic textures, high noise levels, occlusions, or insufficient overlap between images. Given the variable imaging conditions underwater and the complex, intricate structures of coral reefs, such as their tentacles, these challenging scenarios frequently arise, posing substantial demands on the accuracy, robustness, and efficiency of underwater 3D reconstruction technologies.
@@
Despite the impressive performance of these advanced methods in testing environments, their applicability in underwater environments, particularly in coral reef scenes, remains to be further validated. Can these methods truly outperform traditional approaches when applied to coral reef images? While several studies have investigated their potential for high-quality coral reef modeling (Zhong et al. (2023b, 2025, 2024)), a thorough and methodical comparison across methods is still lacking. Addressing this gap is essential for practical applications. Yet, without a comprehensive review and evaluation of both classical and cutting edge methods, researchers and practitioners lack the necessary evidence to select the most effective workflows. This absence of guidance hinders efforts to deploy scalable, reliable surveying solutions at a time when coral ecosystems are under unprecedented threat. This study therefore aims to:
  • systematically review the image based 3D reconstruction algorithms, with a focus on recent technological advancements, and provide a structured synthesis of their development and capabilities.
  • benchmark the performance of various methods on both real and synthetic coral reef datasets, using a comprehensive set of evaluation metrics.
  • provide evidence based recommendations for selecting the optimal solutions of coral reef modeling and discuss current limitations and identify promising directions for future research.
By bridging the divide between photogrammetric computer vision innovations and real world reef applications, our work delivers actionable guidance to support urgent conservation efforts with reliable, high fidelity 3D reconstruction solutions.
The rest of this paper is organized as follows: Section 2 first presents a typical 3D reconstruction workflow for underwater coral images, and then briefly reviews the technical developments addressing specific challenges, including comparative analyses of various methods. Section 3 provides a series of experimental evaluations of the 3D reconstruction solutions discussed in Section 2, performed using various coral reef image datasets. Section 4 further discusses the results presented in Section 3 and explores current challenges and potential future research directions. Section 5 concludes with key findings and their broader implications.
2 3D Reconstruction Workflow and Method Evolution
@@ 2.1 Data collection and preparation
In the first stage, the primary objective is to acquire underwater images using visual sensors such as Digital Single Lens Reflex (DSLR) cameras and action cameras. The captured images should be high-resolution, exhibit sufficient overlap, and collectively cover the entire survey area to enable the application of photogrammetric computer vision techniques for 3D reconstruction of coral reef structures. The spatial resolution of the imagery directly determines the achievable reconstruction accuracy. For example, when monitoring subtle changes in coral morphology, which may be limited to only a few centimeters over the course of a year, it is preferable to achieve millimeter-level resolution. In practice, data collection is typically performed by divers or unmanned vehicles such as Remotely Operated Vehicles (ROVs) or Autonomous Underwater Vehicles (AUVs).
@@ 2.2 Camera pose estimation
Camera pose estimation focuses on determining the accurate positions and orientations of the cameras. Traditional aerial triangulation typically relies on initial estimates of camera poses to achieve accurate results, which can be a limitation in underwater environments that often lack reliable navigation data. In contrast, SfM automatically estimates camera poses without prior information, making it a powerful solution. Its speed, low cost, simplicity, and versatility have led to widespread adoption in 3D reconstruction tasks.
As illustrated in Figure 1, SfM is primarily composed of two main components: correspondence search and reconstruction. The first component aims to detect regions of overlap between input images and match corresponding projections of the same 3D points across overlapping areas. This process involves two critical steps: feature extraction and feature matching. Feature extraction identifies distinctive points (keypoints) in images along with their descriptors, which characterize the surrounding area of each keypoint. These features are extracted using local feature methods designed to be invariant to changes in translation, rotation, scale, and illumination. Subsequently, the matching process seeks to identify corresponding projections of the same points across overlapping images, thereby establishing correspondences between image pairs. Feature matching commonly involves geometrical verification to eliminate potential outliers in the matches (Zhong et al. (2023c)). This verification is typically based on local photometric and geometric constraints. The quality of the correspondence search directly impacts the accuracy and reliability of the reconstruction. Existing solutions may face significant challenges in complex application scenarios, highlighting this as a key research problem.
The second component aims to estimate the image poses and 3D scene points using the correspondences between image pairs, which is achieved through image registration and triangulation in SfM reconstruction. The critical processing step in this phase is bundle adjustment (Triggs et al. (2000)). Given that the extracted feature locations and correspondences may contain errors, and lens distortions can occur, errors can accumulate rapidly, potentially leading to drift in SfM. Bundle adjustment addresses these issues by minimizing the reprojection error—the discrepancy between the projected positions of 3D points in images and the actual locations of detected feature points—thereby improving the accuracy and consistency of the reconstruction. The objective function of bundle adjustment is presented as follow:
@@
where $P_{j}$ represents the pose of an image, $X_{i}$ denotes a 3D scene point, $\pi(\cdot)$ is the function that projects the 3D point onto the image, $x_{j}$ indicates the location of a feature on the image, and $\rho_{ij}(\cdot)$ is a loss function used to potentially down-weight outliers. The notation $\|\cdot\|_{2}$ represents the L2 norm. This optimization process effectively reduces and corrects issues related to noise, mismatches, and initial estimation errors, resulting in refined camera poses and an accurate reconstructed scene structure.
Given the established framework for SfM processes, including pose estimation, triangulation, and bundle adjustment, the correspondence search emerges as the primary challenge, particularly in underwater coral environments. This process becomes unreliable in scenes with chaotic or sparse textures, or under significant variations in viewpoint or lighting conditions, which can hinder SfM reconstruction. Due to the greater impact of the water medium on imaging compared to the air medium, coral reef images captured underwater face challenges such as image degradation, similarity of natural textures, and cluttered background textures. These underwater-specific factors make correspondence search more challenging than in terrestrial applications. Based on a flexible framework for incremental SfM illustrated in Figure 2, the following subsections systematically review three critical technologies within this pipeline: feature extraction, feature matching, and reconstruction.
@@
2.2.1 Feature extraction
SfM relies on local features to align overlapping images by identifying the same scene points. These features must be invariant to radiometric and geometric changes to ensure accurate correspondence between different images. Traditional hand-crafted feature extraction methods typically follow a two-stage pipeline: keypoint detection followed by descriptor computing. The geometric invariance of these descriptors primarily addresses scale invariance and orientation invariance, which have been extensively studied over the past decades. SIFT (Lowe (1999, 2004)) is widely known for its scale and orientation invariance by estimating keypoint scales using gradient histograms. SURF (Bay et al. (2006)) significantly improves the computational efficiency of SIFT through various optimizations, while ORB features (Rublee et al. (2011)) further simplify and accelerate the process. Alcantarilla et al. introduced KAZE features (Alcantarilla et al. (2012)), which leverage non-linear scale space through non-linear diffusion filtering to achieve invariance to rotation, scale, and limited affine transformations. They also proposed an enhanced version, AKAZE (Alcantarilla and Solutions (2013)), which improves processing speed. However, these methods are based on heuristics, which can be difficult to generalize across different scenarios, potentially leading to reconstruction failures.
@@ 2.2.1 Feature extraction
Hand-crafted methods SIFT 1999
Advantages: (1) Rapid computational performance; (2) Simple and comprehensible principles; (3) High accuracy in keypoint localization.
Disadvantages: (1) Significant sensitivity to parameter settings; (2) Performance instability under extreme or challenging conditions.
@@
Deep learning-based methods LIFT 2016
Advantages: (1) Enhanced descriptive capability and robustness; (2) Simple parameter adjustment.
Disadvantages: (1) Greater complexity of the model architecture; (2) Dependence on the datasets; (3) Require more computational power.
@@
2.2.2 Feature matching
After feature extraction, a set of keypoints and their corresponding descriptors are obtained. For SfM reconstruction, accurate feature matching is required to establish correspondences between images of the same scene region. The problem is classically solved by matching a keypoint with its most similar counterpart in another image, specifically identifying the nearest neighbor in descriptor space. However, under conditions such as low image overlap, sparse or overly similar textures, and variations in lighting, descriptors may fail to accurately represent features. This can result in many outliers—matches that are incorrect or significantly different from the expected correspondences. The classical ratio-test (Lowe (2004)) is a basic and straightforward feature matching method, often used in conjunction with mutual nearest neighbor checks for rapid correspondence search. While these simple filters are efficient and widely applied, their performance is limited. They often miss many outliers or incorrectly filter out inliers, which can complicate SfM solutions. To achieve more robust and reliable feature matching, researchers have explored techniques such as local spatial consistency checks and global geometric verification. Among these methods, one of the most representative is AdaLAM (Cavalli et al. (2020)), proposed by Cavalli et al. AdaLAM demonstrates significant robustness and is notable for its ease of deployment and high operational efficiency. While AdaLAM integrates several ideas from previous research, it remains based on geometrical assumptions, such as local affine consistency, which may limit its applicability in certain scenarios—especially when dealing with no-planar surfaces. To address this, it employs adaptive strategies to relax these assumptions, enhancing its generalization across different domains. Compared to general methods, AdaLAM not only significantly reduces outliers but also increases the number of inliers.
@@ 2.2.3 SfM reconstruction
2.3 Dense surface reconstruction
Although SfM reconstruction produces 3D point clouds through triangulation, the resulting point clouds are typically too sparse to represent the detailed geometric information of the observed coral reef areas. This limitation is particularly pronounced in underwater coral reef environments, where fine-scale structures such as small coral branches or tentacles are of ecological significance. To address this, dense surface reconstruction techniques are employed to recover high-resolution representations of the scene. This limitation necessitates the use of dense surface reconstruction techniques to recover the finer morphological details essential for accurate monitoring and analysis.
Dense surface reconstruction refers to the process of estimating 3D surface geometry by integrating information across multiple image views. Contemporary methods for dense reconstruction employ a range of approaches, but they all share a common principle of generating a dense representation of the 3D scene, such as a dense point cloud or mesh. This process depends on the geometric relationships and viewpoint consistency among multi-view images for inference and optimization. Consequently, it necessitates the intrinsic and extrinsic camera parameters, which can be derived from SfM reconstruction. The quality of the dense surface reconstruction directly influences downstream tasks such as measurement, scene analysis, and digital modeling.
In the context of coral reef observation, dense surface reconstruction offers the ability to model fine morphological details, facilitating precise structural analysis. However, coral reefs present unique challenges due to their complex and diverse forms. Features such as densely packed coral polyps, overhangs, and branched structures often lead to occlusions and self-similarity, which can confound reconstruction algorithms. Environmental factors such as uneven lighting, water turbidity, and suspended particles further degrade image quality, introducing noise and reducing feature consistency across views. Moreover, structural changes in coral reefs can occur on the scale of mere centimeters over annual timescales, demanding reconstruction methods with exceptionally high spatial accuracy. Given these challenges, there is an urgent need for high-quality and efficient dense reconstruction solutions. As illustrated in Figure 3, current dense surface reconstruction methods can be broadly categorized into four categories: traditional multi-view stereo (MVS) methods, deep learning-based MVS methods, methods based on neural radiance fields (NeRFs), and methods based on Gaussian Splatting (GS). Each category offers distinct advantages and limitations in terms of reconstruction quality and computational efficiency. Representative approaches within these categories are summarized in Table 2 and will be discussed in detail in the following subsections.